OpenStudy (anonymous):

You are installing a new spark plug in your car, and the manual specifies that it be tightened to a torque that has a magnitude of 47.1 N·m. Using the data in the drawing below (L = 0.297 m and θ = 49.6°), determine the magnitude F of the force that you must exert on the wrench. You are installing a new spark plug in your car, and the manual specifies that it be tightened to a torque that has a magnitude of 47.1 N·m. Using the data in the drawing below (L = 0.297 m and θ = 49.6°), determine the magnitude F of the force that you must exert on the wrench. @Mathematics

OpenStudy (matt101):

As you probably know, the equation for torque is: (I'm assuming L is the lever arm) T= L*F*sinθ 47.1=0.297*F*sin(49.6) --> F = 421.4 N

OpenStudy (anonymous):

this the diagram...pls the anser u gave aint right

OpenStudy (anonymous):

try resolve it using this diagram:

OpenStudy (matt101):

Need a username and password for that...my answer was probably wrong because L isn't the lever arm, so yeah a diagram would be helpful haha

OpenStudy (anonymous):

yeah click the link i gave u

OpenStudy (matt101):

That link needs you to have log in info...you're probably logged in which it why it takes you to an actual page instead of the log in screen.

OpenStudy (anonymous):

wow wuld u need my password? i can give it to u just for tonight

OpenStudy (anonymous):